Consider the LR circuit below. Suppose we close switch S1. When the current in the circuit has reached a steady state value,just parts c, d, and g please
Answer #1

c)Just after S2 is closed, current through R2 is zero because the induced emf produced in the coil opposes applied current. Hence current can easily pass through S2 (It is not a resistive path).

d) Just after S2 is closed, current flows through the path S1R1S2V.

Hence current through S2 i=V/R1=48/56=0.857A

g) Current through S2 long time after S2 is closed =0.857A

This is because at the steady state even the coil does not produce induced emf, resistance R2 opposes current. Hence current easily flows through S2 as there is no resistance in this path.